### v3

# When Geoscience Meets Foundation Models: Towards General Geoscience Artificial Intelligence System

Abstract
Geoscience foundation models (GFMs) represent a revolutionary approach within Earth sciences to integrate massive cross-disciplinary data for improved simulation and understanding of Earth system dynamics. As a data-centric artificial intelligence paradigm, GFMs extract valuable insights from petabytes of both structured and unstructured data. Their versatility in task specification, diverse inputs and outputs, and multi-modal knowledge representation enable a comprehensive analysis that surpasses the capabilities of individual data sources. Critically, the scalability and generalizability of GFMs empower them to address a wide array of prediction, simulation, and decision tasks related to the intricate interactions among Earth system components. By unraveling the causal mechanisms underlying observed patterns and changes, GFMs contribute to advancing our knowledge of the Earth system and its responses to various drivers and perturbations. Collaboration between domain experts and computer scientists plays a pivotal role in fostering innovations in these invaluable tools for understanding the past, present, and future of our planet. Moreover, we introduce recent advances in GFMs including key technologies for constructing GFMs, large language models, large vision models, and large vision-language models, especially remote sensing applications. However, challenges remain in validation and verification, scalability, interpretability, knowledge representation, and addressing social bias. Going forward, the key lies in enhancing model integration, resolution, accuracy, and equity through interdisciplinary teamwork. Despite current limitations, GFMs hold great promise for providing critical insights into pressing issues including climate change, natural hazards, and sustainability through their ability to explore multiple scenarios and quantify uncertainties. Their continued evolution toward integrated, data-driven modeling holds paradigm-shifting potential for Earth science.
Index Terms:
Foundation model, Geoscience artificial intelligence, Data-centric paradigm, Causal inference, Multimodal learning
I Introduction
AS Earth science enters the era of big data, artificial intelligence (AI) offers great potential not only for solving classic geoscientific problems, but also for enhancing understanding of Earth’s complex, interactive, and multiscale processes [1, 2]. The availability of massive volumes of Earth system data, already exceeding dozens of petabytes with hundreds of terabytes transmitted daily, has led to the widespread utilization of AI, including machine learning and deep learning, in data-driven Earth system science [3, 4]. For example, deep learning has been successfully applied to identify extreme weather patterns [5] to build competitive weather prediction models ranging from precipitation nowcasting [6, 7] to medium-range weather seasonal forecasts [8, 9, 10, 11], and to predict climate phenomena such as ENSO [12] or monsoon onsets [13]. Deep learning models can also automatically extract patterns and features from vast amounts of satellite and aerial imagery, which enables a wide range of applications, including environmental monitoring, land use and land cover classification, urban planning, disaster response, and climate change analysis [14, 15, 16, 17]. Moreover, generative models such as generative adversarial networks or diffusion models promise to be very valuable for bias correcting and upsampling Earth system model fields [18, 19, 20, 21]. However, fully utilizing the explosive growth and diversity of Earth system data presents two major challenges in the coming years: (i) conventional AI models have limited ability to extract knowledge from the data deluge, and (ii) supervised models cannot readily adapt to varying tasks through one training, limiting generalization. To leverage the massive amounts of Earth system data, larger models are urgently needed for data-driven Earth system science.
Foundation models—large pre-trained models representing the newest generation of AI—are trained on vast, diverse datasets and can be fine-tuned for various downstream tasks [22]. These models achieve outstanding performance across wide ranges of tasks, from answering textual queries [23] to describing images [24] or playing video games [25]. The adaptability of foundation models is a key feature that sets them apart from earlier AI models. Earlier AI models were designed for specific singular tasks and lacked the capacity to perform well on diverse tasks outside of their original purpose. In contrast, the extensive pre-training of foundation models on large and varied datasets allows them to learn rich representations of language and visual data.
Foundation models offer unprecedented capabilities, driven by the expansion of datasets, the scaling up of model size, and advancements in model architectures. For instance, as of 2023, the language model ChatGPT-4 and Claude-3 have incorporated vision, speech, and knowledge graphs, astounding users with their ability to generate text closely resembling human writing and create images and computer code from various prompts [26, 27, 23]. Moreover, many contemporary foundation models can process inputs from and produce outputs across multiple domains. Early research has explored medical foundation models capable of working with both images and text [28, 29, 30, 31]. As certain capabilities only emerge at larger model sizes, it remains difficult to predict the future accomplishments of even bigger models.
Although there have been preliminary efforts to develop geoscience foundation models (GFMs) [32, 33, 21], this shift has not yet become widespread in geoscience AI, owing to the persisting challenges in accessing expansive, diverse Earth system datasets and the complexity of the Earth system, as well as the recency of this advancement. As a result, geoscience AI models predominantly continue to be developed using a task-specific approach. For instance, a weather prediction model may be trained solely on historical climate records from a specific region, likely requiring substantial data collection efforts. This model would only be able to forecast climate change for that region over a few days, and would be unable to predict rainfall-induced shallow landslides [34]. Moreover, a remote sensing model trained on satellite images from arid regions might fail when applied to classify wetlands in Southeast Asia. The difference in vegetation, water bodies, and terrain means the model may inaccurately identify land features, leading to incorrect analyses of such ecologically sensitive areas (Aslam et al., 24; Chen et al., 2024) [35, 36]. Such narrow, task-specific models yield inflexible systems, constrained to performing predefined tasks based on the training dataset and labels. Currently, these models typically cannot adapt to new tasks or different data distributions without complete retraining. Although the task-specific approach persists as the dominant paradigm in geoscience AI, emerging foundation models may enable more flexible, multifaceted systems given sufficient accessible Earth system data.
Recent advancements in foundation models research have the potential to disrupt the prevailing task-specific paradigm in the geoscience by enabling a more versatile and adaptable approach to data analysis and prediction. Key developments include the emergence of multimodal architectures and self-supervised learning techniques that do not require explicit labels, such as language modeling and contrastive learning [37, 38]. Additionally, the advent of in-context learning capabilities allows foundation models to rapidly adapt to new tasks and data distributions. These innovations may transform how geoscientists process and interpret data, leading to enhanced accuracy and impact in critical research areas like sustainability, climate change, extreme weather events, natural hazards, and environmental monitoring [39]. As foundation models continue to evolve and increase in scale, their potential to revolutionize geoscience research and applications is immense. Adoption of these adaptable, multifaceted foundation models in lieu of narrow, static task-specific models may yield new insights and capabilities in the geoscience.
Inspired directly by foundation models outside geoscience, we identify three crucial capabilities that distinguish GFMs from conventional geoscience AI models. Firstly, adapting a GFM for new task will become much easier. These models will be able to address novel problems simply by receiving a textual explanation of the new task (flexible task specification), without requiring full retraining. Secondly, GFMs can accept varied combinations of multimodal data as input and generate diverse multimodal outputs. This enables processing of diverse data types including images, text, laboratory results, or any combination thereof. This flexible interactivity contrasts with inflexible multi-modal models that rely on predefined modalities for inputs and outputs. For example, such models must always take in a fixed set of images, text, and lab results together. Thirdly, GFMs will incorporate formal representations of geoscience system knowledge, empowering analysis of novel tasks and communication of outputs using precise professional language. Overall, GFMs are distinguished by their flexible task specification, versatile multimodal capabilities, and formal geoscience knowledge representations.
II Key advantages of geoscience foundation models
Figure 1: Overview of the foundation model. a. The development of AI. Machine learning pays its attention on “how” to define the features of data and make decision using the learning algorithms, while deep learning learns features automatically by designing specific architectures. Foundation models focus on functionalities such as in-context learning emerge by using large pre-trained models. b. The characteristics of foundation models. Foundation models have five characteristics, large data, large pre-trained models, large computing force, self-training, and wide applications. c. The paradigm shift from model-centric AI to data-centric AI.
II-A Foundation models: a new paradigm
Machine learning and deep learning have been widely applied for solving problems in Earth system science [3, 1, 40, 41]. As presented in Figure 1a, machine learning focuses on “how” to define the features of data and make decisions using learning algorithms, while deep learning infers “features” automatically by designing specific architectures. Foundation models focus on functionalities such as in-context learning by using large pre-trained models. They rely on massive, diverse data, large pre-trained models, and self-supervised training, yet exhibit even larger computing demands.
II-A1 Data-centric AI
Traditional machine learning methods focus on developing models and algorithms for predefined datasets and in particular deep learning involves unsupervised learning tasks. However, real-world applications involve working with noisy, imperfect data, and uncertain labels. Rather than treating datasets as fixed, performance can be substantially improved by refining the data itself. This has given rise to the emerging field of Data-Centric AI (DCAI), which aims to systematically engineer better datasets as a means to improve model performance. Whereas skilled data scientists have long manually curated datasets in an ad hoc fashion based on intuition, DCAI seeks to establish more formal and objective principles and techniques for dataset optimization. This represents a paradigm shift from perceiving data as a static input to models, to co-developing datasets and models in an integrated fashion. Further research in DCAI may uncover new methods for improved data collection, automated data cleaning, labeling, augmentation, and representation learning that make the overall machine learning process more robust and efficient.
Drawing inspiration from foundational models beyond geoscience, we have identified three distinct features that set GFMs apart from traditional geoscience machine and deep learning models (Figure 1b). Firstly, employing GFMs for novel tasks will be as straightforward as articulating the tasks in natural language processing, such as English or other languages. GFMs will possess the capability to address previously unencountered challenges through dynamic task specification, enabling them to comprehend and tackle new tasks without the need for retraining. Secondly, GFMs can accept inputs and generate outputs utilizing various combinations of data modalities. This flexibility allows for the incorporation of different types of data, such as videos, images, sound, text, laboratory results, or any combination thereof, into the model. This flexible interactivity capability contrasts with the constraints of more stringent traditional multimodal models. Thirdly, GFMs will employ a formal representation of geoscientific knowledge, enabling it to engage in reasoning processes for previously unseen tasks and provide interpretations using geoscientifically accurate language.
II-A2 Scalability and generalizability
As illustrated in Figure 1c, foundation models include large amounts of data, large models, requiring large computing capacity, self-training, all enabling wide ranges of applications, exhibiting excellent scalability and generalizability. The scalability is enabled along three key dimensions: data, model size, and computing power, and generalizability refers to the ability of a model to perform well on new tasks beyond the data it was trained on.
Scalability
First, foundation models are trained on massive datasets, frequently comprising hundreds of billions of data samples. The models benefit from this vast data, allowing them to learn very broad patterns. Second, the models themselves are scaled up to huge sizes, frequently with billions or even trillions of parameters. By increasing model capacity, more nuanced representations can be learned from the extensive data. Finally, substantial computing resources are leveraged to train and run these massive models, with thousands of GPUs/TPUs used in parallel computing grids. Together, this facilitates unprecedented scalability, allowing foundation models to achieve capabilities not previously possible with limited data and traditional deep learning models. Their scalability underpins their versatility across a broad array of downstream tasks.
Generalizability
Foundation models exhibit strong generalizability due to their training methodology. Specifically, foundation models are pretrained on diverse datasets in a self-supervised manner, learning broad representations of language, vision, robotics, or other modalities. This differs from traditional supervised learning, where models are tailored to specific labeled datasets. The self-supervised pretraining endows foundation models with extensive world knowledge and versatility. Subsequently, the pretrained models can be adapted and fine-tuned for myriad downstream tasks using relatively little labeled data. Even unseen tasks can be potentially tackled in a zero-shot setting without any downstream data. Their flexibility and wide applicability highlight the generalizability of foundation models across domains. Their generalizability originates from the self-supervised pretraining and enables adaption to new distributions and tasks.
II-A3 Dynamic interactions
Conventional AI models are restricted to one or several data types as input, and they are even sensitive to the data shape, i.e., 2D or 3D data can often not be processed within the same model due to the limitation of the convolutional operation. Moreover, AI models are always designed for specific tasks, such as image classification or object detection. These limitations constraint the applications of traditional machine and deep learning models, thus, GFMs need to be further developed to enable the handling of multi-modal, multi-type inputs and outputs and so executing flexible tasks.
Multi-modal inputs and outputs
GFMs aim to represent and understand the physical processes and properties of the Earth’s interior, crust, surface and atmosphere. These models take in multiple types of input data (multimodal inputs) including seismic data, gravity data, heat flow measurements, laboratory rheology data, weather data or land-surface data e.g. from remote sensing. Each data type provides constraints on different aspects of the Earth system. Models can generate multiple types of output (multimodal outputs) including synthetic seismic waves, temperature/pressure fields, surface displacement maps, etc. Handling multi-modal data allows the models to take advantage of complementary constraints from different geophysical observables. This can reduce non-uniqueness and ambiguity compared to using a single data type alone.
Flexible task
GFMs aim to represent complex natural processes in Earth system like mantle convection, crustal deformation, groundwater flow, ocean and atmospheric dynamics, or vegetation. These processes interact in non-linear ways across multiple spatial and temporal scales. GFMs can be taught to flexibly specify new tasks without requiring models to be retrained. For instance, a weather prediction model can be realized by training on historical climate data, and this model can also be helpful for the forecast of rainfall-induced shallow landslides.
Causal inference
The integration of GFMs into the exploration of Earth’s processes and phenomena heralds a transformative shift in our ability to discern and understand the intricate causal relationships inherent in geoscience knowledge [42]. By leveraging data from diverse sources and modalities—ranging from satellite imagery and sensor data to historical climate records and geological samples—GFMs offer a novel approach to modeling the Earth system that is both holistic and nuanced. This comprehensive data integration enables GFMs to capture the complexity and interconnectedness of the Earth system’s components. For instance, by analyzing patterns across atmospheric data, oceanic temperatures, and land use changes, GFMs can help researchers identify underlying causal mechanisms driving climate change, such as the feedback loops between greenhouse gas emissions and global warming [43, 44]. Similarly, in the context of natural disasters, GFMs can elucidate the cascading effects of a single event, like how an earthquake can trigger tsunamis, which in turn affect coastal erosion and sediment transport. The potential of GFMs to elucidate causal relationships in geoscience is grounded in their ability to process and analyze vast datasets that would be unmanageable for traditional analytical methods. Through advanced machine learning techniques, GFMs can detect patterns and correlations within this data, proposing hypotheses about causal links that can be further tested through empirical research or simulation models. This not only accelerates the discovery of new knowledge but also enhances our ability to predict future Earth system changes and inform policy and decision-making. Moreover, the flexibility of GFMs to incorporate data from new sources and update their models in real-time allows for the continuous refinement of our understanding of causal relationships as new evidence emerges. This adaptability is crucial in geoscience, where ongoing changes in the Earth system and technological advancements in data collection methods constantly evolve the landscape of available information.
II-B Geoscience domain knowledge
As shown in Figure 2, geoscience encompasses diverse domains spanning the physical makeup and processes of the Earth. Core geoscience domains include geology, geography, geophysics, hydrology, oceanography, meteorology, and planetary science. Each domain has accumulated extensive observations, data, and models of a wide range of Earth system components. For instance, geology involves studying the composition, structure, and history of the Earth’s crust and interior, accumulating knowledge on rock types, plate tectonics, mineral resources, and geologic time [45]. Meanwhile, meteorology focuses on the atmosphere and atmospheric phenomena like weather patterns, clouds, and precipitation. Key meteorological knowledge includes air masses, fronts, forecasting models, and climate dynamics [46]. Moreover, geography tends to explore the spatial patterns and processes of natural phenomena, such as climate change, natural disasters, and biodiversity, as well as human aspects like urbanization, cultural landscapes, and economic development. Geographers require a diverse set of knowledge, including understanding physical geography, human geography, and technical skills in geographic information systems (GIS), remote sensing, and spatial analysis [47]. Beyond domain-specific knowledge, geosciences also rely on integrative theories and principles regarding Earth systems interactions, such as the carbon cycle, the hydrologic cycle, or interactions between highly nonlinear components of the Earth system that may exhibit abrupt regime transitions [48]. Together, robust domain knowledge across the geosciences enables understanding the complex, interconnected processes that shape the Earth and its environments. Accumulating, codifying, and connecting this knowledge is an ongoing endeavor as geoscience domains continue to expand their theoretical and empirical foundations through observation, experimentation, and modeling.
Figure 2: Overview of the GFM. By leveraging techniques like self-supervised learning, the GFMs undergo training using multimodal and multidimensional geoscience data models. For enhanced versatility, the data schema can be coupled with language inputs, such as text, voice, or video data. Furthermore, the GFMs require access to diverse sources of geoscientific knowledge to facilitate geoscientific inference tasks, unlocking a plethora of functionalities applicable to downstream applications. Once trained, the GFMs can dynamically perform user-specified tasks in real-time. To accomplish this, the model can retrieve contextual information from geological mapping manuscripts, knowledge graphs, databases, and other relevant sources, capitalizing on formal geoscience knowledge to deduce solutions for previously unseen tasks.
II-C Potential Applications
Figure 3: Illustration of potential applications of Geoscience foundation model (GFM). On the left side, Earth science is depicted as encompassing four prominent disciplines: solid Earth science, atmospheric science, planetary science, and marine science. On the right side, various observation methods currently utilized in Earth science are represented, including space-based, air-based, ground-based, and ocean-based data collection. These observation methods provide multi-modal and multi-dimensional data that contribute to the field of Earth science. In the middle lies the GFM, which has the capability to integrate, correlate, and analyze the diverse geoscience data obtained from different observation means. By extracting higher-level features and representations from the data, GFMs enable enhanced understanding and insights. Once trained, this model can dynamically perform user-specified tasks in real-time, providing valuable support for Earth science research and applications.
The geoscience field encompasses a diverse range of data types, each fulfilling a distinct role in research and analysis. These data types are pivotal in unraveling the mysteries of our planet and informing various scientific inquiries. As illustrated in Figure 3, the diverse array of observational modalities (space-air-ground-ocean) employed in geoscience research each provide unique affordances for characterizing the Earth system. Space-based data collection provides a global perspective, offering comprehensive coverage of the Earth’s surface including vegetation and ecosystems, the atmosphere and oceans. Air-based data collection serves as a bridge between space and terrestrial observations, enabling localized, high-resolution data acquisition. Ground-based data collection involves direct, in-situ measurements, enhancing our comprehension of geographical or geological occurrences. Additionally, ocean-based data collection plays a crucial role in evaluating marine and climatic conditions.
Moreover, GFMs offer significant potential for causal discovery form Apart from multimodal data, considering the diverse observational modalities such as space, air, ground, and ocean. Numerous research questions in Earth sciences are inherently causal, requiring robust analyses to establish the causal relationships among variables from multimodal data [42]. By integrating data from various sources and modalities, GFMs enable researchers to uncover causal links between different variables and phenomena, allowing for a more comprehensive and holistic understanding of Earth processes, as it takes into account the interconnectedness of different components of the Earth system. By unraveling the causal mechanisms underlying observed patterns and changes, GFMs contribute to advancing our knowledge of the Earth system and its responses to various drivers and perturbations.
II-C1 Space
AI models have demonstrated their capability to integrate the characteristics of the biosphere and ecosystems on Earth, thereby providing valuable assistance in the search and assessment of signs of life on other planets [49]. We anticipate the development of GFMs that can integrate multimodal and high-dimensional data from the field of planetary science, which aims to obtain more comprehensive information and in-depth insights. GFMs promise to serve as powerful tools in planetary science by facilitating the fusion, correlation, and analysis of data from various modalities. For example, by simultaneously analyzing planetary surface image data, remote sensing data, and radar data, GFMs can facilitate the identification of dependencies and connections among different data sources, thereby enabling more accurate extraction of geomorphic features, identification of geological structures, and prediction of crustal activities. Moreover, with its standardized and adaptable framework for data integration, analysis, and visualization, GFMs are poised to play a pivotal role in unveiling fresh insights and discoveries concerning celestial bodies across our solar system. As space missions become more ambitious, GFMs’ ability to facilitate efficient and collaborative data processing will become increasingly important in advancing our comprehension of planetary processes and environments.
II-C2 Air
AI models have gained widespread recognition and utilization in the field of weather forecasting [8, 9, 10, 11], with promising advances also regarding extreme weather events such as extreme rainfall [19], long-term droughts, heavy rainfall, typhoons, and strong winds [8]. To enhance the accuracy and timeliness of weather forecasts, there is a need for advanced GFMs that can effectively integrate large volumes of multi-modal meteorological observation data and utilize sophisticated modeling techniques. Such a model would enable meteorologists and forecasters to make more accurate predictions and stakeholders to make more informed evidence-based decisions, leading to improved preparedness, and enhanced understanding of meteorological phenomena. For instance, weather forecasting relies on data from multiple sources, such as satellites [50], radars, weather stations, and numerical models. GFMs can serve as a unified platform to integrate and fuse these diverse data streams, providing a comprehensive view of atmospheric conditions. Moreover, the recent success of generative machine learning models such as generative adversarial networks, to bias correct and downscale climate models fields [19] suggests that GFM can also play a crucial role for paving the way for improved assessments of climate change impacts. By combining data from different sources, forecasters can make more informed and accurate predictions. GFMs will play a crucial role by providing a standardized semantic representation for weather data, ensuring consistency in data formats, units, and structures. This standardized approach facilitates data sharing, collaboration, and interoperability between different meteorological agencies and research institutions. Moreover, weather forecasting is increasingly reliant on comprehending the intricate interactions between the atmosphere, ocean, land, and cryosphere. GFMs can integrate Earth system models, enabling forecasters to consider complex feedback mechanisms and enhance the accuracy of long-range forecasts. Consequently, GFMs can significantly improve the accuracy and speed of weather forecasting in the future by leveraging their ability to process vast amounts of multimodal data, enable advanced data analytics, and facilitate interdisciplinary collaborations.
Moreover, the integration of foundation models within the field of remote sensing image segmentation represents a paradigm shift, significantly elevating the precision and efficiency of categorizing complex imagery into distinct, meaningful classes such as vegetation cover, water bodies, and urban infrastructure [51, 52]. Leveraging sophisticated deep learning architectures, these foundation models adeptly navigate the intrinsic variability and intricacy of spatial patterns captured in satellite and aerial imagery. This proficiency is paramount for a spectrum of critical applications, encompassing environmental monitoring (e.g., tracking deforestation, urban expansion), disaster response mechanisms, and agricultural optimization, among others. Foundation models distinguish themselves by their capacity to directly learn high-dimensional feature representations from extensive datasets of annotated images, thereby obviating the need for labor-intensive manual feature extraction. This capability not only streamlines the segmentation workflow but also significantly enhances the adaptability and scalability of remote sensing analysis. Consequently, the adoption of foundation models is not only merely enhancing the operational efficacy of image segmentation tasks, but also concurrently expanding the horizons for innovative research and applications in remote sensing, fostering a more nuanced comprehension of terrestrial phenomena and the intricate interplay between natural and anthropogenic landscapes.
II-C3 Ground
AI models have made significant advancements in the field of geophysics, contributing to various areas such as earthquake prediction and early warning systems, geophysical exploration, and geological risk assessment [1]. Seismic disasters often have significant impacts on society when they occur, yet seismologists currently lack the ability to accurately predict the exact timing of earthquakes. We anticipate the development of comprehensive GFMs that can greatly enhance our understanding of earthquake processes, improve the accuracy of earthquake predictions, and provide valuable information and decision support for earthquake disaster management and risk assessment. For example, earthquake prediction relies on various types of multidimensional seismic data, such as seismic waveforms, seismic records, and underground geological data. Advanced GFMs excel at processing high-dimensional data and integrating multiple modalities of seismic information, thereby enabling more comprehensive and accurate earthquake prediction outcomes. Furthermore, it is essential to acknowledge that although GFMs can offer substantial contributions to earthquake prediction, this field remains exceptionally challenging due to the intricate nature of seismic processes. Continuous research endeavors, advancements in data collection methodologies, and collaborative initiatives are pivotal in further enhancing the accuracy and efficiency of earthquake prediction. Nevertheless, GFMs hold tremendous potential in revolutionizing earthquake forecasting and mitigating the potential impact of seismic events on communities and critical infrastructure.
II-C4 Ocean
AI models enable holistic monitoring, modeling, and forecasting of conditions in the near-surface ocean through multi-sensor data analysis, digital twin simulations, and long-term observations at multi-scale [53]. However, compared with the near-surface ocean, the intermediate ocean at a depth from 100 to 1000 m is confronted with a series of challenges primarily stemming from the scarceness of observational data, which has emerged as a significant impediment constraining the progressive evolution of marine science and data-driven modelling in oceanography. We anticipate the development of a GFM featuring real-time, immersive, and highly sensitive dynamic interactive visualization capabilities for three-dimensional oceanic data. This model holds the promise of gradually unraveling the enigma surrounding the intermediate and deep oceans, thereby ushering future marine science into a transformative new phase of exploration and understanding. For instance, advanced GFMs have the potential to catalyze a transformation in midocean remote sensing technology, offering the capability for autonomous, long-term monitoring of poorly explored mid-ocean regions. This can be achieved through intelligent fusion and analysis of diverse subsurface, environmental, and biological data streams from distributed, next-generation ocean observation platforms, thereby revolutionizing the way mid-ocean regions are studied and understood. Moreover, GFMs play a transformative role within the framework of Deep Blue AI [54], facilitating advanced data analysis, predictive modeling, and well-informed decision-making to elevate the quality of marine science research, exploration, and conservation endeavors, thereby helping to unraveling the enigma of one of Earth’s least-explored yet profoundly significant frontiers.
III Recent Advances in Geoscience Foundation Models
With the applications of large foundation models for general purposes, the explorations of foundation models for remote sensing-based Geoscience tasks have drawn a lot of attention from the research community. In this subsection, we will review the recent advances in GFMs, including the key techniques of constructing GFMs, the summary of existing foundation models from the view of large language models (LLMs), large vision models (LVMs), and large language-vision models (LLVMs). Table I summarizes the existing foundation models, including data type, pre-training, adaptation, applications, and key features.
III-A Key Techniques of Constructing Geoscience Foundation Models
As shown in Figure 4, pre-training and adaptation are two critical stages in the development and application of foundation models, where the former involves training a model on a large, diverse dataset to learn general-purpose representations, and the latter involves fine-tuning the pre-trained model on a specific task or dataset. The pre-training stage is particularly important for foundation models, as it enables the model to learn general-purpose representations that can be adapted to a wide range of tasks. The adaptation stage is equally important, as it allows the model to be fine-tuned for specific tasks or datasets, thereby improving its performance on those tasks.
Figure 4: Key Techniques of Constructing Geoscience Foundation Models, Pre-training (a-d) and Fine-tuning (e-g).. a. Supervised Pre-training, training with large labeled data. b. Generative Pre-training, training by reconstructing data. c. Contrastive Pre-training, training by contrastive learning using similarity. d. Hybrid Pre-training, training with labeled and unlabeled data. e. Model Fine-tuning, fine-tuning GFMs using domain specific datasets. f. Prompt Engineering, fine-tuning GFMs using prompts. g. Prompt Tuning, fine-tuning GFMs using learned prompts.
III-A1 Pre-training
Pre-training involves training a foundation model on a large, diverse dataset. The primary goal is to learn a broad range of patterns, knowledge, and skills that can be universally applied across various tasks. Pre-training methods for GFM include supervised pre-training, self-supervised pre-training, and hybrid pre-training, where self-supervised pre-training involves generative and contrastive pre-training [55].
Supervised pre-training
Supervised pre-training involves training a model on a large, labeled dataset. The model learns to predict the labels of the data, and the weights of the model are updated to minimize the prediction error. Supervised pre-training is particularly useful when there is a large, labeled dataset available for pre-training.
Self-supervised pre-training
Self-supervised pre-training involves training a model on a large, unlabeled dataset. The model learns to predict the missing parts of the data, and the weights of the model are updated to minimize the prediction error. Self-supervised pre-training is particularly useful when there is a large, unlabeled dataset available for pre-training. Generative pre-training trains a model to generate data that is similar to the input data by generating or reconstructing input data, while contrastive pre-training trains a model to distinguish between similar and dissimilar data.
Hybrid pre-training
Hybrid pre-training involves combining supervised and self-supervised pre-training. The model is first pre-trained on a large, labeled dataset using supervised pre-training, and then pre-trained on a large, unlabeled dataset using self-supervised pre-training. Hybrid pre-training is particularly useful when there is a large, labeled dataset available for pre-training, and a large, unlabeled dataset available for pre-training.
III-A2 Adaptation
Pre-trained GFM models can be adapted to specific tasks or datasets through model fine-tuning, prompt tuning, and prompt engineering.
Model fine-tuning
Model fine-tuning involves updating the weights of the pre-trained model on a specific task or dataset. The weights of the model are updated to minimize the prediction error on the specific task or dataset. Model fine-tuning is particularly useful when there is a small, labeled dataset available for adaptation.
Prompt tuning
Prompt tuning involves updating the prompt of the pre-trained model on a specific task or dataset. The prompt of the model is updated to minimize the prediction error on the specific task or dataset. Prompt tuning is particularly useful when there is a small, labeled dataset available for adaptation.
Prompt engineering
Prompt engineering involves designing a prompt for the pre-trained model to perform a specific task or dataset. The prompt is designed to guide the model to generate the desired output for the specific task or dataset. Prompt engineering is particularly useful when there is a small, labeled dataset available for adaptation.
TABLE I: Summary of Existing Foundation Models
Domain Ref. Data Type Pre-training Adaptation Applications Key Features
LLMs
GeoBERT [56]
Geoscientific records
Generative
Question answering and summarization
Retrained on geoscientific records
BERT-E [57] Earth science-related articles Generative - Keyword classification Transfer learning from SciBERT [58]
K2 [21]
Geo-text, geoscience instructions correlations
Model fine-tuning
Geoscience tasks
Geoscience tasks
CnGeoPLM [59] Chinese geological corpus Generative - Geological NLP tasks A Chinese pre-trained LVM
GeoBERTSegmenter [60]
Unregistered geological terms
Generative
Word segmentation
Chinese word segmentation model
OceanGPT [61] Ocean domain instruction data Generative - Oceans science tasks The first-ever LLM in the ocean domain
Lawley et al. [62]
Geo-scientific reports and publications
Generative
NLP tasks
Observation-based data analysis
Lawley et al. [63] Geoscientific text data Generative - Extract semantic relationships New knowledge extraction
GeoGalactica [64]
geoscience text corpus
Generative
Model fine-tuning
Question answering & knowledge discovery
Largest language model
LVMs
RingMo [14], Wang et al. [15], Cha et al. [65]
Remote sensing images
Generative
Model fine-tuning
Remote sensing image analysis
RingMo-Sense [66] Remote sensing images Generative Model fine-tuning Spatiotemporal prediction Video Swin Transformer
RingMo-Lite [67]
Remote sensing images
Generative
RS image interpretation
Optimized for edge devices
RingMo-SAM [52] Remote sensing images Generative Model fine-tuning Remote sensing image segmentation Effective across various datasets
SAMRS [68], RSPrompter [36]
Remote sensing images
Prompt engineering
Remote sensing image segmentation
Segment Anything Model (SAM)
Ma et al. [69] Remote sensing images Generative Model fine-tuning Remote sensing image segmentation Segment Anything Model (SAM)
Geo-SAM [70]
Remote sensing images
Model fine-tuning
Infrastructure segmentation
Segment Anything Model (SAM)
OFA-Net [71] Remote sensing images - Model fine-tuning Infrastructure segmentation Segment Anything Model (SAM)
CSP [72]
Remote sensing images
Contrastive
Model fine-tuning
Satellite image classification
Satellite
Prithvi [16] Satellite imagery Generative Model fine-tuning Earth observation tasks open source
ClimaX [73], W-MAE [74]
Grid-based meteorological data
Generative
Model fine-tuning
Climate change prediction
Climate and weather modeling
Fourcastnet [9], GraphCast [10], Fengwu [11], Pangu-Weather [8] Grid-based meteorological data Generative - Weather forecasting Climate and weather modeling
SkySense [75]
Multi-modal Remote Sensing Imagery dataset
Contrastive
Model fine-tuning
Earth image analysis
State-of-the-art performance
LVLMs
RSVQA[76]
Images and text
Generative
Model fine-tuning
Satellite image analysis
Satellite image analysis
LiT-4-RSVQA [77] Images and text Generative Model fine-tuning Vision question answering Lightweight
Chappuis et al. [78]
Exploration of language biases in RSVQA
GeoChat [79] Multimodal geoscience data - Prompt tuning Geospatial data analysis Robust zero-shot performance
EarthGPT [80]
Image-text pairs
Instruction tuning
Universal RS image comprehension
Open-set reasoning tasks.
SkyeyeGPT [81] RS instruction-following dataset - Instruction tuning Image-level and region-level tasks Comparable results with GPT-4V
Mall et al. [82]
Multimodal geospatial data
Generative
Model fine-tuning
Geospatial data analysis
Satellite images
RS-CLIP [83] Image–text datasets Contrastive Model fine-tuning Remote sensing scene classification Zero-shot remote sensing scene classification
Zhang et al. [84]
Images and text
Evaluation of LVLMs in Earth observation tasks
Datasets[85] Image-text datasets - - Remote sensing RSICap and RSIEval datasets for VLMs
III-B Large Language Models
Large language models (LLMs) have been widely used in many applications since BERT [86] was introduced. In the geoscience domain, LLMs have been used for various tasks, including geoscience question answering, query-based summarization, keyword classification, and geoscientific text data analysis.
Several LLMs for geoscience have been developed based on BERT [86], SciBERT [58], and LLaMA [87]. For example, Denli et al. [56] retrained a pre-trained BERT model with our 20M internal geoscientific records, referring to GeoBERT. They fine-tuned the GeoBERT model for several tasks including geoscience question answering and query-based summarization. The work [57] introduced BERT-E, an Earth science-specific language model built using transfer learning from SciBERT [58], which shows improved performance in keyword classification and highlights the need for robust benchmarks in evaluating such domain-specific language models, with potential applications across various scientific fields. The authors in [21] introduced K2, the first LLM specifically for geoscience, along with a new instruction tuning dataset, GeoSignal, and a benchmark, GeoBench, to adapt and evaluate LLMs in geoscience, demonstrating its effectiveness in understanding and utilizing geoscience knowledge with open-source data and model checkpoints.
For Chinese geological texts, the authors in [59] developed a large-scale Chinese geological corpus (GeoCorpus) and a Chinese pretrained language model (CnGeoPLM) for geological domain, demonstrating that CnGeoPLM significantly outperforms the generic BERT model in geological NLP tasks like entity recognition, relationship extraction, and entity clustering. Wei et al. [60] introduced GeoBERTSegmenter, a GeoBERT-based Chinese word segmentation model tailored for geological texts, combining BERT, bidirectional recurrent neural network, and conditional random field to effectively address the segmentation of unregistered geological terms and common words, demonstrating superior performance in precision, recall, and F1 score compared to other methods.
Despite general geoscience tasks, LLMs have been used in specific geoscience domains. For example, the authors in [61] presented OceanGPT, the first-ever Large Language Model specialized for ocean science, developed with a novel framework, DoInstruct, for generating domain-specific instruction data and evaluated using OceanBench, a new oceanography benchmark, demonstrating enhanced expertise and preliminary embodied intelligence capabilities in ocean science tasks.
Additionally, the authors in [62] described the development and validation of new geoscience-specific language models trained on public geoscientific reports and publications, which outperform generic models in various NLP tasks and demonstrate the potential of using NLP to extract meaningful semantic relationships and latent knowledge from the vast volume of geoscience literature. [63] described the application of natural language processing (NLP) to geoscientific text data to extract meaningful semantic relationships and predict mineral potential, demonstrating how geoscience language models can transform unstructured text into actionable insights to streamline mineral exploration for critical materials. [64] presented the development of GeoGalactica, the largest language model specialized in geoscience, which is trained on a vast geoscience text corpus and fine-tuned with instruction-tuning data, highlighting the model’s comprehensive training process and its open-source contribution to the field.
In summary, the development of LLMs for geoscience has been an active area of research, with several models developed and evaluated for various geoscience tasks. These models have demonstrated the potential of using LLMs to extract meaningful insights from geoscience text data and to improve the performance of geoscience NLP tasks.
III-C Large Vision Models
Large vision models (LVMs) have been widely used in remote sensing and geoscience for various tasks, including remote sensing image analysis, satellite image classification, and climate and weather modeling. Most of the LVMs are based on the transformer architecture, which is effective in handling large-scale remote sensing data. For example, the authors in [14] introduced RingMo, a remote sensing (RS) foundation model framework using generative self-supervised learning, trained on a large-scale dataset of two million RS images to address the domain gap and generalization issues in RS image interpretation, showing state-of-the-art performance across multiple datasets and tasks, advocating for the adoption of generative self-supervised learning in RS research. The work [15] introduced LVMs specifically tailored for remote sensing tasks, featuring a new rotated varied-size window attention mechanism in vision transformers, demonstrating superior performance in detection, classification, and segmentation tasks with improved computational efficiency and data transferability. Cha et al. [65] discussed the impact of increasing the number of model parameters on the performance of foundation models in remote sensing tasks like rotated object detection and semantic segmentation, demonstrating that larger models improve performance and data efficiency, achieving state-of-the-art results on several benchmark datasets. The authors in [66] presented RingMo-Sense, a remote sensing foundation model for spatiotemporal prediction, utilizing spatial, temporal, and spatiotemporal triple-branch prediction networks with parameter-sharing and progressive joint training strategies, validated on an RS spatiotemporal dataset and demonstrating competitive performance across six downstream tasks. [67] introduced RingMo-lite, a lightweight CNN-Transformer hybrid network optimized for RS image interpretation on edge devices, effectively handling high- and low-frequency spectral components and reducing computational resources while maintaining high accuracy and state-of-the-art performance in various RS tasks, soon to be integrated into the MindSpore computing platform.
Moreover, several LVMS for geoscience are based on the Segment Anything Model (SAM), which is a transformer-based model that is effective in handling remote sensing data. For example, Yan et al. [52] introduced RingMo-SAM, a foundation model for multimodal remote-sensing image segmentation, capable of segmenting and identifying object categories in both optical and SAR data, using a large-scale dataset and a novel category-decoupling mask decoder and prompt encoder, demonstrating effectiveness across various datasets. The authors in [68] introduced SAMRS, a large-scale remote sensing segmentation dataset created using the Segment Anything Model (SAM) and existing RS object detection datasets, containing over 105,090 images and 1.6 million instances, with a comprehensive analysis and preliminary experiments demonstrating its utility for various segmentation and detection tasks, addressing challenges of limited training data. [36] presented the development of RSPrompter, an automated instance segmentation approach for remote sensing images based on the Segment Anything Model (SAM), which leverages prompt learning to improve SAM’s ability to generate semantically meaningful segmentation results, validated by extensive experiments on various datasets, with the code publicly available. Ma et al. [69] described a novel framework utilizing the Segment Anything Model (SAM) for semantic segmentation in remote sensing, introducing SAM-Generated Object (SGO) and SAM-Generated Boundary (SGB) concepts along with novel object and boundary loss components, demonstrating improved performance on two datasets, with the source code available online. The authors in [70] introduced Geographical SAM (GeoSAM), an advanced SAM-based framework enhanced with dense and sparse visual prompts from zero-shot learning and a pre-trained CNN, which significantly outperforms existing methods in segmenting mobility infrastructure in geographical images, marking a significant advancement in the use of foundation models for this purpose. The work [71] exploited OFA-Net (One-For-All Network), a versatile method using a single Transformer backbone for multiple data modalities and spatial resolutions in remote sensing, pre-trained on a multi-modal dataset, showing promising performance across 12 different downstream tasks, moving towards a unified foundation model in Earth vision.
For satellite image classification, the authors in [72] introduced CSP, a contrastive pre-trained model for satellite image classification, which outperforms existing methods in satellite image classification tasks, demonstrating the effectiveness of contrastive pre-training in remote sensing tasks. Similarly, [16] introduced Prithvi, a transformer-based geospatial foundational model pre-trained on over 1TB of multispectral satellite imagery, demonstrating its effectiveness in various Earth observation tasks through efficient pre-training and fine-tuning, with significant improvements over existing models and the ability to maintain accuracy with limited labeled data, made publicly available as an open-source contribution.
The authors in [73] presented ClimaX, a flexible and generalizable deep learning model for weather and climate science, extending the Transformer architecture to handle heterogeneous datasets and various spatio-temporal scales, demonstrating superior performance in weather forecasting and climate projections compared to other data-driven models, with the source code available online. The work [74] introduced W-MAE, a novel weather and climate model based on the Transformer architecture, which outperforms existing models in weather forecasting and climate change prediction, demonstrating the potential of large vision models in weather and climate science. Similarly, Fourcastnet [9], GraphCast [10], Fengwu [11], and Pangu-Weather [8] are all large vision models specifically tailored for weather forecasting and climate modeling, demonstrating the potential of large vision models in weather and climate science. Guo et al. citeguo2023skysense exploited SkySense, a large vision model for Earth image analysis, which outperforms existing models in Earth image analysis tasks, demonstrating the potential of large vision models in Earth image analysis.
To summarize, LVMs have been widely used in remote sensing and geoscience for various tasks, including remote sensing image analysis, satellite image classification, and climate and weather modeling. These models have demonstrated the potential of using large vision models to improve the performance of remote sensing and geoscience tasks.
III-D Large Vision-Language Models
Large vision-language models (LVLMs) are increasingly being used in the fields of remote sensing and geoscience. LVLMs are often utilized for visual question answering. The authors in [76] first described how Remote Sensing Visual Question Answering (RSVQA) leverages advancements in natural language processing and computer vision to analyze satellite images with language processing, comparing the effectiveness of different language models, particularly focusing on the benefits of fine-tuning in enhancing performance. The work [77] presented LiT-4-RSVQA, a lightweight transformer-based architecture for efficient and accurate visual question answering in remote sensing, which reduces computational demands while maintaining accuracy, as demonstrated by its performance on a VQA benchmark dataset. Chappuis et al. [78] outlined how Remote Sensing Visual Question Answering (RSVQA) facilitates human-machine interaction with overhead imagery using natural language, emphasizing the need to address language biases in RSVQA through a comprehensive analysis involving visual blind models, adversarial testing, and dataset analysis, revealing more severe biases in remote sensing compared to standard VQA due to dataset specifics.
Additionally, LVLMs have been used for remote sensing image understanding. Kuckreja et al. [79] introduced GeoChat, the first remote sensing Vision-Language Model with multitask conversational capabilities, designed to overcome the limitations of general-domain VLMs in handling high-resolution, category-diverse remote sensing imagery. GeoChat excels in a variety of RS tasks and addresses the challenge of a lack of domain-specific datasets by creating a novel RS multimodal instruction-following dataset. The authors in [80] outlined EarthGPT, a large vision-language model for Earth observation data, which outperforms existing models in Earth observation tasks, demonstrating the potential of large vision-language models in Earth observation. Zhan et al. [81] introduced SkyeyeGPT, a large vision-language model for remote sensing, which outperforms existing models in remote sensing tasks, demonstrating the potential of large vision-language models in remote sensing. The work [85] discussed the development of a high-quality Remote Sensing Image Captioning dataset (RSICap) and a benchmark evaluation dataset (RSIEval) to advance VLMs in remote sensing, addressing the lack of comprehensive image-text datasets for training and effectively evaluating these models in the domain. Mall et al. [82] introduced a method training LVLMs for remote-sensing images without textual annotations by aligning an image encoder for remote sensing images with CLIP’s image encoder using paired internet and satellite images, enabling zero-shot, open-vocabulary image classification, retrieval, segmentation, and visual question answering, outperforming supervised models on several tasks. Similarly, Li et al. [83] introduced RS-CLIP, an LVLM for remote sensing scene classification that uses contrastive vision-language supervision and pseudo-labeling techniques, achieving high zero-shot classification accuracy on four benchmark datasets by learning semantic-aware visual representations and applying a curriculum learning strategy.
Zhang et al. [84] discussed the evaluation of LVLMs on Earth observation data through a comprehensive benchmark focusing on scene understanding, localization, counting, and change detection tasks, revealing that while these models excel in tasks like location understanding and captioning, their spatial reasoning limitations affect performance in object localization and counting. The authors in [85] discussed the development of a high-quality Remote Sensing Image Captioning dataset (RSICap) and a benchmark evaluation dataset (RSIEval) to advance LVLMs in remote sensing, addressing the lack of comprehensive image-text datasets for training and effectively evaluating these models in the domain.
LVLMs have been widely used in remote sensing and geoscience for various tasks, including visual question answering and remote sensing image understanding. These models have demonstrated the potential of using large vision-language models to improve the performance of remote sensing and geoscience tasks.
IV Challenges of foundation models in Geoscience AI
IV-A Validation and Verification
GFMs possess the unique capacity to process complex inputs and generate likewise complex outputs. This makes validating their correctness exceptionally challenging, if not impossible. Unlike conventional geoscience AI models designed for specific tasks with supervised learning approaches validated within predefined use case parameters, GFMs promise to perform novel tasks specified by end-users without further training. Therefore, anticipating all possible failure modes for such broadly capable systems a prior becomes categorically more difficult. Consequently, developers and regulators bear the responsibility of transparently delineating how GFMs have been tested and which use cases they have been approved to operate within. Comprehensive testing methodologies and use case definitions will be critical for ensuring proper functioning and, hence, trust.
Many geoscience processes operate over prolonged timescales, including climate change [8, 7], ice sheet dynamics, or tectonic motions, necessitating consistent Earth observations spanning decades to centuries or even-in that case from proxy data at paleoclimate time scales, to sufficiently characterize them. Testing GFMs requires amassing illustrative cases over comparably long time periods to encapsulate rare events statistically. Even within a human lifespan, acquiring robust statistical samples poses challenges. Furthermore, when foundation models fail at predicting major events, responding systematically to identify specific weaknesses proves more constructive than overreacting to isolated failures. Often, targeted model refinements emerge from rigorous post-event analyses, contingent on collecting appropriate data preceding the event. Adopting a measured, evidence-based approach for incrementally improving models through their occasional failures will promote greater confidence and utility over the long term.
IV-B Multi-Source Data
Foundation models are trained on very large datasets and typically have billions of parameters. The computational requirements for training are massive, needing thousands of GPUs/TPUs and costing substantial energy resources. Moreover, extensive computing for training and inference has huge environmental impacts without conscientious mitigation efforts, especially carbon emissions. Costs and computing requirements concentrate progress in few organizations, risking monopoly power over impactful technologies.
To enhance our comprehension of the intricate diversity within geoscience data, a systematic classification encompassing space, air, ground, and ocean realms is imperative. This classification scheme offers a refined perspective on the multifaceted nature of these data types. The ensuing enumeration delineates prevalent data categories prevalent within the field: (i) Satellite and remote sensing data: Geoscientists extensively leverage satellite imagery and remote sensing data, which encompass optical, infrared, radar, and lidar datasets [88]. These sources provide invaluable insights into land cover, vegetation distribution, topographical variations, geological formations, and dynamic environmental changes. (ii) Geological and geophysical data: Geoscientists gather and analyze a plethora of geological and geophysical data, including seismic readings, gravity, and magnetic measurements, well logs, core samples, and geological maps [89]. These datasets are fundamental in decoding subsurface structures, discerning rock properties, and comprehending geological processes. (iii) Climate and weather data: Comprising parameters like temperature, precipitation patterns, humidity levels, solar radiation, wind dynamics, and atmospheric pressure, climate and weather data [90], which play a pivotal role in unraveling climate change phenomena, enhancing weather forecasting accuracy, and illuminating intricate interactions within the atmosphere, land, cryosphere and oceans. (iv) Geospatial data: Geospatial data integration, facilitated through Geographic Information Systems (GIS), harmonizes diverse Earth science datasets such as georeferenced maps, spatial databases, and geographic coordinates [91]. This integration empowers spatial analysis, visual representation, and predictive modeling. (v) Environmental monitoring data: This category encapsulates essential parameters like air and water quality, soil composition, and pollutant concentrations [92]. Environmental monitoring data is instrumental in scrutinizing shifts in environmental conditions, conducting comprehensive pollution assessments, and devising sustainable resource management strategies. (vi) Time-series data: Time-series observations, spanning climate records, seismic datasets, flow measurements, and glacier motion data, provide a chronicle of temporal changes [42]. Time-series analysis offers a means to identify patterns, track trends, and detect anomalies within these dynamic datasets. (vii) Text and scientific literature data: Geoscientists extensively tap into textual resources extracted from scientific literature, research publications, field reports, and scholarly notes [1]. These textual reservoirs not only furnish valuable insights but also serve as crucial references for guiding and informing ongoing scientific investigations.
IV-C Interpretability
Foundation models like GPT-4 are difficult to interpret because of their scale and complexity. With billions of parameters, it is challenging to explain why they generate a particular output. The inner workings seem opaque compared to smaller, more specialized AI systems. On the other hand, foundation models have some degree of interpretability. Techniques like attention mechanisms provide some visibility into what input the model is focusing on. There are also methods to analyze feature importance and conduct ablation studies to understand how different components of the model contribute to its outputs. The natural language capabilities of models like GPT-3 also lend themselves to certain interrogation techniques. Overall, there are challenges in interpreting large foundation models, but there are also opportunities. Interpretability is an active area of research, and new techniques are being developed. There is always a spectrum of interpretability, where no model is perfectly transparent or perfectly opaque. Researchers continue to explore methods to open the “black box” of foundation models while also recognizing their limitations. There are reasonable arguments on multiple sides of this complex issue.
IV-D Domain Knowledge
The complexity of geoscience domain knowledge arises from several intrinsic challenges. Firstly, it spans a broad spectrum of disciplines, including geology, geophysics, meteorology climate science, oceanography, and seismology, among others. Given that much of the Earth’s interior and history remains inaccessible for direct observation, integrating knowledge across these diverse fields is inherently complex. Secondly, geoscience operates from sub-daily to geological timescales, encompassing approximately 4.5 billion years of Earth system evolution. This extensive temporal scope means that distilling even fundamental knowledge becomes an immensely challenging task. Thirdly, the Earth system is characterized by intricate interconnections between the geosphere, hydrosphere, cryosphere, atmosphere, and biosphere. Understanding the dynamics of these interconnected mechanisms requires a holistic, interdisciplinary perspective. Furthermore, geoscience phenomena occur across a vast range of scales, from microscopic to planetary, making direct observation and sampling challenging. Lastly, the field of geoscience is continuously evolving, with new discoveries frequently expanding, revising, and refining our understanding of the Earth’s processes. Consequently, summaries and overviews can quickly become outdated as knowledge advances. This dynamic nature of geoscience knowledge further adds to the complexity of defining a comprehensive domain understanding.
IV-E Transparency and Trust
Previous research and reactions within scientific communities have revealed that the implementation of AI in geosciences raises concerns among geoscientists. Specifically, GFMs are often regarded as black boxes, presenting challenges in comprehending the rationale behind their predictions or explanations [93]. Geoscientists who emphasize transparency in their research may have reservations about using GFMs. Moreover, geoscientists working in data-scarce fields or those facing limited access to pertinent datasets may encounter challenges in the applicability and effectiveness of GFMs. When the available dataset inadequately represents specific scenarios or harbors harmful correlations, the risk of bias in test results during the training of GFMs becomes more pronounced. Ensuring the absence of undesirable biases in these unprecedentedly large and complex training datasets poses a formidable task, with the potential for social biases to amplify alongside the scale of the model. Additionally, geoscientists may be concerned about over-reliance on AI, which could result in a detrimental erosion of human expertise and critical thinking within the decision-making process.
GFMs necessitate rigorous verification across diverse scenarios prior to their promotion and application, guaranteeing their robust performance in critical and specialized contexts. Moreover, ongoing auditing and governance are imperative even after the models are deployed, considering the emergence of new challenges as they encounter novel tasks and settings. Notably, incentivizing the AI community through prize-winning competitions can serve as a catalyst for meticulous scrutiny of GFMs. For instance, participants could be rewarded for identifying indicators that engender harmful content or expose potential failure modes. Swiftly identifying and rectifying biases must remain a paramount objective for developers, suppliers, and regulators alike.
V Conclusion
GFMs hold immense promise for a significant advancement in modeling and prediction of Earth’s systems dynamics, thus enhancing our understanding of these intricate processes. Operating as a groundbreaking paradigm, GFMs inherit the idea of data-centric AI by leveraging large datasets and computational power to uncover concealed insights. Through flexible task specification, multi-modal inputs and outputs, and multi-modal knowledge, these GFMs models offer comprehensive insights that are infeasible through individual data sources or methodologies. Most importantly, the scalability and generalizability of GFMs make them uniquely capable of tackling a wide range of prediction, simulation, and decision-making challenges related to the Earth system dynamics, encompassing the highly nonlinear components and their complex multi-scale interactions. Going forward, fostering collaboration across various Earth science disciplines, in conjunction with computer scientists and applied mathematicians, will be pivotal in ushering in the next generation of GFMs. With their unique ability to explore multiple scenarios and quantify uncertainties, these models promise to provide critical insights into pressing issues such as climate change, natural hazard assessment, and sustainable resource management. Recent advancements in GFMs have demonstrated their potential to revolutionize the field of Earth system modeling, especially in the context of climate and weather forecasting, remote sensing, and geoscience. Despite prevailing challenges in validation and verification, scalability, interpretability, domain knowledge, and addressing social biases, GFMs have become an indispensable part of analyzing the past, monitoring the present, and planning for the future of the Earth.
Acknowledgment
The authors would like to express their sincere gratitude and acknowledgment to Professor Gregory C. Beroza from Stanford University for his invaluable opinion on this paper.